Our bodies absorb oxygen (O2) through our lungs, then distribute this oxygen to our muscles and organs. We have found that the average person has 0.0423 g O2 in their blood. 

(a) If the average person has 4.70 L of blood in their bodies, what is the concentration of oxygen (in Molarity) in the average person’s blood? (assume that blood is just water)

 

 

​(b) A person can go into shock if they lose too much blood. What mass (g) of oxygen would be lost if a person lost 0.78 L of blood?



1
Expert's answer
2021-02-22T05:55:50-0500

(a) Molar mass of O2 = 32 g/mol

Average mass of O2 in blood = 0.0423 g

Moles of O2 = "\dfrac{0.0423}{32}=1.321875\times10^{-3} mol"

Volume of Blood = 4.70 L

"\therefore" Molarity = "\dfrac{1.321875\times10^{-3}}{4.70}\dfrac{mol}{L}"

Molarity = "2.8125\times10^{-4}mol/L"


(b) 1L of blood contains = "2.8125\times10^{-4}" moles of O2

0.78 L of blood contains = "2.8125\times10^{-4}\times0.78" moles of O2

= "2.19375\times10^{-4}" moles of O2

Mass of O2 lost = "2.19375\times10^{-4}\times32" g

= "7.02\times10^{-3}" g